Five grams of sugar are added to a 45-gram serving of a breakfast cereal that is 10% sugar.What is the percent concentration of sugar in the resulting mixture?

Answer by stanbon(75887)

Five grams of sugar are added to a 45-gram serving of a breakfast cereal that is 10% sugar.What is the percent concentration of sugar in the resulting mixture?
---
sugar + sugar = sugar
1*5 + 0.10*45 = x(5+45)
-----
0.10*45 = 50x
x = (9/10)0.1
x = 9/100
x = 9%
=============
Cheers,
Stan H.
==============